# HP-BERT: A framework for longitudinal study of Hinduphobia on social media via LLMs

### v1

@@ Abstract
During the COVID-19 pandemic, community tensions intensified, fuelling Hinduphobic sentiments and discrimination against individuals of Hindu descent within India and worldwide. Large language models (LLMs) have become prominent in natural language processing (NLP) tasks and social media analysis, enabling longitudinal studies of platforms like X (formerly Twitter) for specific issues during COVID-19. We present an abuse detection and sentiment analysis framework that offers a longitudinal analysis of Hinduphobia on X (Twitter) during and after the COVID-19 pandemic. This framework assesses the prevalence and intensity of Hinduphobic discourse, capturing elements such as derogatory jokes and racist remarks through sentiment analysis and abuse detection from pre-trained and fine-tuned LLMs. Additionally, we curate and publish a ”Hinduphobic COVID-19 X (Twitter) Dataset” of 8,000 tweets annotated for Hinduphobic abuse detection, which is used to fine-tune a BERT model, resulting in the development of the Hinduphobic BERT (HP-BERT) model. We then further fine-tune HP-BERT using the SenWave dataset for multi-label sentiment analysis. Our study encompasses approximately 27.4 million tweets from six countries, including Australia, Brazil, India, Indonesia, Japan, and the United Kingdom. Our findings reveal a strong correlation between spikes in COVID-19 cases and surges in Hinduphobic rhetoric, highlighting how political narratives, misinformation, and targeted jokes contributed to communal polarisation. These insights provide valuable guidance for developing strategies to mitigate communal tensions in future crises, both locally and globally. We advocate implementing automated monitoring and removal of such content on social media to curb divisive discourse.
keywords:
Hinduphobia , COVID-19 , X (Twitter) , Large Language Models
journal: Nuclear
@@
[inst1]organization=Department of Data Science and Artificial Intelligence,addressline=IIIT, Naya Raipur, postcode=Chhattisgarh, country=India
[…]
[inst2]organization=Transitional Artificial Intelligence Research Group, School of Mathematics and Statistics, addressline=UNSW Sydney, postcode=Sydney, country=Australia
[…]
[inst3]organization=Centre for Hindu Studies, addressline=Pingala Institute, postcode=Sydney, country=Australia
1 Introduction
@@
A global pandemic emerged from the severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) [1], which resulted in the coronavirus disease 2019 (COVID-19)[2] with a profound impact on social and economic activities worldwide [3, 4, 5]. One of the most significant aspects of this pandemic has been how social media has influenced public attitudes and behaviours related to the pandemic [6]. The COVID-19 pandemic sparked a global public health emergency and led to a rise in online discussions that exacerbated existing social divisions and sparked new conflicts within communities worldwide [7]. In India’s case, a country known for its diverse cultural landscape and inclusive principles, the pandemic exacerbated existing socio-political divisions, particularly concerning communal interactions [8]. Although efforts have focused on addressing anti-Islamic sentiments and behaviours [9, 10, 11, 12], there is another significant but less explored issue: the escalation of Hinduphobia [13, 14] during the COVID-19 crisis. Furthermore, the COVID-19 pandemic has also brought to light the issue of Sinophobia [15, 16, 17], but there has been little attention given by researchers to the rise of Hinduphobia.
@@
Hinduism is the world’s oldest practising religion [18] with a population of more than 1.1 billion followers worldwide who mostly reside in India. Hinduism encompasses some of the oldest philosophical (Upanishads and Bhagavad Gita) [19, 20], epic and literary texts (Ramayana and Mahabharata) [21, 22]. Hinduism has contributed to modern science and technology with the invention of the binary and decimal number system [23] (also known as Hindu and Arabic numerals) [24] and systematic approach to scientific investigation using the philosophy of pramana (evidence) [25, 26]. Hinduism is based on non-violent principles that promote vegetarianism [27], with India having the largest vegetarian population in the world with the majority of Hindus being partially vegetarian (specific days of vegetarian diet) [27]. Hinduism urges taking a violent stand in self-defence [28] and the fight for dharma (ethics) as promoted in the Bhagavad Gita [29]. The term Hinduphobia refers to hate speech and racial and cultural slurs about Hindu religious and cultural practises [13]. Although limited, few studies have investigated the nature of Hinduphobia in television and popular culture [13], which became more prominent with the rise of the Internet and social media [14].
@@
Hinduism in India withstood its grounds with a history of violence and discrimination given 800 years of foreign rule, with invasions leading to temple desecration [30]. Thousands of Hindu temples have been destroyed desecrated and converted into mosques and churches in history due to foreign rule with religious and political intervention [30, 31]. This has been continuing with Hinduphobia in social media and political discourses, especially in countries where Hindus have been a minority, such as Pakistan and Bangladesh [32, 33]. In a majority Hindu India, Hindupbois and anti-Hindu political rhetoric led to the genocide and exodus of Hindus [34] from the Kashmir valley region in the 1980s [35]. Therefore, Hinduphobia is not a recent phenomenon, it has more than a thousand years of history given the invasions and foreign rule of India [36]. Furthermore, Hindupbobic rhetoric has been one of the reasons for forces conversion of Hindu women in Pakistan [37] and gradual changes in ownership and conversion of Hindu temples [38]. This has been an issue not just in India but in countries with a Hindu history such as Indonesia [39].
@@
The Kumbh Mela which is a major Hindu festival in India was associated with a high number of COVID-19 cases during the Delta variant outbreak in April 2022 [40, 41, 42]. This association, driven by a political media narrative, led to negative portrayals of the Hindu community on social media platforms including Twitter, Reddit, Instagram, and Facebook. These negative portrayals resulted in trends with hashtags such as #HinduVirus and #CoronaHindus, contributing to Hindubhobic statements and hate speech against the Hindu community. Previous studies have examined Hinduphobia in contexts of historical events, political dynamics, academic biases, interreligious relations, diaspora experiences, and cultural practices [13, 43]. Furthermore, there has been an upsurge in hate crimes against Hindus, as demonstrated by the cases of Khaniya Lal, Umesh Kohle, and Nupur Sharma [44]. These incidents highlight the escalating hostility towards Hindus during the pandemic and the need for religious tolerance in society. The COVID-19 pandemic has exacerbated communal tensions in India, which motivates our study to systematically study the rise of Hinduphobic sentiments connected to COVID-19 in India and the rest of the world.
@@
In recent years, LLMs have emerged as powerful tools for analysing and understanding public sentiment on social media. LLMs, such as (GPT-3, PaLM, and GPT-4) [45, 46] can process vast amounts of textual data, revealing underlying patterns, sentiments, and discourse dynamics that shape public perception and societal narratives [47]. These models can be based on pre-training from a large data corpus and achieve strong performance accuracy in tasks such as decision-making, predictions, text generation and text summarizations [48]. LLMs can detect nuanced language patterns and emerging trends, enabling researchers and policymakers to respond more effectively to rising social and political tensions [49] taking into account security and privacy [48]. Addressing the root causes of discrimination and fostering social cohesion is crucial; however, applying LLMs in this context has a lot of challenges. Data privacy, ethical considerations, and the potential for algorithmic bias must be carefully managed to ensure the deployment of LLMs for societal well-being [50]. LLM must limit biases and maintain fairness, particularly when dealing with sensitive topics such as communal sentiments and hate speech [51].
Sentiment and semantic analysis have become crucial components of NLP applications, particularly for analysing vast amounts of unstructured textual data from social media platforms [52, 53]. These methods involve understanding the emotional tone (sentiment) and the meaning (semantics) of textual content to extract insights regarding public opinion, trends, and biases [54]. LLMs have significantly improved the accuracy and scalability of these tasks by leveraging massive datasets and deep model architectures. LLMs can capture complex linguistic patterns and contextual nuances previously challenging to NLP. One of the earliest LLMs, the Bidirectional Encoder Representations from Transformers (BERT) model [55], has been widely adopted for sentiment analysis tasks [56]. Several customised BERT models have been developed, including news recommender BERT [57], fake news BERT [58], legal-BERT [59, 60] and medical-BERT [61] for respective domains. The Multilingual-BERT (mBERT) [62] has been pre-trained on monolingual corpora in 104 languages, which has been designed to detect hate speech across multiple languages. DistilBERT [63] is a variant of BERT optimised for resource-limited applications. Moreover, advanced variants such as Ro-BERT-a [64] have proven effective for nuanced hate speech detection across diverse settings. In the case of detection of abuse and violence in NLP applications, BERT has been designed for various contexts and categories of hate speech. HateBERT [65] is trained on RAL-E, a large-scale dataset of Reddit comments from communities banned for being offensive, abusive, and hateful. These examples of customised BERT models and their variants demonstrate the adaptability and power of LLMs in addressing the complexities of hate speech detection in specific contexts, showcasing their impact in real-world NLP applications.
Sentiment and semantic analysis have been used to study various domains using LLMs, including social media and text analysis. Wang and Chandra [17] used BERT-based sentiment analysis and reported that surges in COVID-19 cases correlated with spikes in Sinophobic sentiments. Similarly, Chandra et al. [66] used BERT-based sentiment analysis to study trends in The Guardian newspaper coverage of COVID-19, revealing a predominance of negative sentiments. In another study, Chandra and Kulkarni [67] utilised semantic and sentiment analysis to evaluate the Bhagavad Gita Sanskrit-English translations, comparing verse-by-verse consistency across selected translators. In the field of machine translation, Wang et al. [68] evaluated Google Translate for its accuracy in translating Mandarin Chinese, applying sentiment and semantic analysis to assess translation precision. Similarly, Shukla et al. [69] analysed Sanskrit-to-English translations of the Bhagavad Gita, uncovering inconsistencies in translating philosophical and metaphorical terms of translation by Google Translate. Furthermore, Chandra and Saini [70] modelled the United States 2020 presidential elections using BERT-based sentiment analysis, providing insights into sentiment trends on social media before the elections. Chandra and Krishna [71] compared LSTM and BERT models to track public sentiment during the rise of COVID-19 cases in India. These studies collectively highlight the role of BERT-based models and LLMs in analysing sentiments across social and news media, illustrating their versatility and potential for understanding public opinion.
@@
In this study, we present the ”Hinduphobic COVID-19 X (Twitter) Dataset, the first dataset of its kind, with over 8,000 tweets collected from November 2019 to December 2022 during the COVID-19 pandemic. It has 2000 manually labelled tweets and the rest employs GPT-3.5 turbo API. The dataset is labelled into two categories: Hinduphobic (negative) and positive/neutral, with all tweets related to Hinduism. This dataset is publicly available via Kaggle [^1] https://www.kaggle.com/datasets/ashutoshsingh22102/hinduphobic-covid-19-x-twitter-dataset-india. We use the Hinduphobia dataset to fine-tune a BERT (HP-BERT) model for classifying Hinduphobic tweets. Through a multi-stage fine-tuning strategy [76, 77, 78], we further train this model on the ”SenWave dataset” [79] to enable sentiment analysis. Subsequently, the model is applied to the ”Global COVID-19 Twitter dataset” [80], which includes tweets from six countries including Australia, Brazil, India, Indonesia, Japan, and the United Kingdom capturing a longitudinal perspective on Hinduphobic sentiment trends across these regions during and post-COVID-19 pandemic. We also analyze sentiment polarity scores to gain deeper insights into how Hinduphobic sentiments evolved and varied geographically. Additionally, we use HateBERT [65] to examine toxicity and abusive language in tweets in the ”Global COVID-19 X (Twitter) Dataset” [80], offering a further understanding of the propagation of Hinduphobia and the behavioural patterns of users spreading such content. We present a refined BERT (HP-BERT) model specifically designed for detecting Hinduphobia, with adaptations to handle Hinglish (Hindi and English) data from India [81], which includes content compiled for abuse and misogyny analysis [75]. This model effectively detects Hinduphobic content and performs sentiment analysis on identified tweets, providing nuanced insights into the emotional tone and context surrounding Hinduphobic discourse.
@@
The rest of the paper is organised as follows: Section 2 reviews the literature relevant to this study. Section 3 describes the datasets used, including the data collection and preprocessing processes, along with details of the training and testing of our BERT (HP-BERT) model. Section 4 introduces the framework for our approach. Section 5 presents the study’s findings, discussing the performance of the HP-BERT model. Finally, Section 6 provides key insights, and Section 7 concludes the paper with future research directions.
2 Related Work
@@
The study of online activity drew significant attention from the research community due to the unique set of challenges faced during the COVID-19 pandemic, with concerns such as anti-vaccination and misinformation [82, 83, 84]. The substantially increased social media activity provided researchers with a wealth of data to analyze public sentiment and discourse. This surge in online interactions also exacerbated stress in the society [85, 86], as individuals grappled with the rapid spread of misinformation, social isolation, and heightened fears about the pandemic. Chen et al. [87] curated a multilingual coronavirus Twitter dataset with over 123 million tweets for analysing the pandemic’s online discourse, reflecting major events and their impacts. Tahmasbi et al. [88] conducted one of the first studies on Sinophobic behaviour during the pandemic, analysing over 222 million tweets to reveal the rise in anti-Chinese sentiment.
@@
Cho et al. [89] focused on privacy concerns related to COVID-19 contact tracing apps such as TraceTogether, exploring cryptographic protocols and privacy-preserving techniques. The study highlighted the balance between privacy protection and public health benefits, emphasising effective anonymisation and additional privacy measures. Ferrara et al. [90] investigated the influence of X (Twitter) bots during the early COVID-19 outbreak, particularly in spreading political conspiracies and misinformation. Their curated dataset revealed how bots amplified divisive content during global emergencies. Kouzy et al. [91] examined COVID-19 misinformation on Twitter, identifying its rapid spread and sources. They found misinformation often originated from informal groups rather than verified accounts, underscoring the need for robust monitoring and fact-checking mechanisms. Tahmasbi et al. [88] analysed Sinophobic content on social media platforms post-COVID-19, using word2vec models to trace the evolution of anti-Chinese sentiment. They highlighted the exacerbation of racial hostility during the pandemic across different online forums.
@@
Figure 1: Data collection and processing workflow from initial data extraction via Twitter API to manually labelling data points. This is followed by automatic labelling of the remaining tweets using GPT-3.5 Turbo using human-in-the-loop strategy [95].
Despite the extensive research on various aspects of the COVID-19 pandemic, there remains a gap in quantitative analysis regarding Hinduphobic sentiments during this period. This paper aims to fill this gap by examining the prevalence and impact of Hinduphobic discourse on social media platforms amidst the COVID-19 pandemic.
3 Methods
3.1 Datasets

### v2

# HP-BERT: A framework for longitudinal study of Hinduphobia on social media via language models

@@ Abstract
During the COVID-19 pandemic, community tensions intensified, contributing to discriminatory sentiments against various religious groups, including Hindu communities. Recent advances in language models have shown promise for social media analysis with potential for longitudinal studies of social media platforms, such as X (Twitter). We present a computational framework for analyzing anti-Hindu sentiment (Hinduphobia) during the COVID-19 period, introducing an abuse detection and sentiment analysis approach for longitudinal analysis on X. We curate and release a ”Hinduphobic COVID-19 XDataset” containing 8,000 annotated and manually verified tweets. We then develop the Hinduphobic BERT (HP-BERT) model using this dataset and achieve 94.72% accuracy, outperforming baseline Transformer-based language models. The model incorporates multi-label sentiment analysis capabilities through additional fine-tuning. Our analysis encompasses approximately 27.4 million tweets from six countries, including Australia, Brazil, India, Indonesia, Japan, and the United Kingdom. Statistical analysis reveals moderate correlations (r = 0.312-0.428) between COVID-19 case increases and Hinduphobic content volume, highlighting how pandemic-related stress may contribute to discriminatory discourse. This study provides evidence of social media-based religious discrimination during a COVID-19 crisis.
Index Terms:
Hinduphobia, COVID-19, X (Twitter), Language Models, Hate Speech Detection, Religious Discrimination, BERT
@@
=-15pt
I Introduction
@@
A global pandemic emerged from the severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) [1], which resulted in the coronavirus disease 2019 (COVID-19)[2] with a profound impact on social and economic activities worldwide [3, 4, 5]. One of the most significant aspects of this pandemic has been how social media has influenced public attitudes and behaviours related to the pandemic [6]. The COVID-19 pandemic sparked a global public health emergency, leading to a rise in online discussions that exacerbated existing social divisions and sparked new conflicts within communities worldwide [7]. In India’s case, a country known for its diverse cultural landscape and inclusive principles, the pandemic exacerbated existing socio-political divisions, particularly concerning communal interactions [8]. Although efforts have focused on addressing anti-Islamic sentiments and behaviours [9, 10, 11, 12], there is another significant but less explored issue: the escalation of Hinduphobia [13, 14] during the COVID-19 crisis. Furthermore, the COVID-19 pandemic has also brought to light the issue of Sinophobia [15, 16, 17], but computational analysis of Hinduphobia during global crises remains critically underexplored, representing a significant gap in hate speech detection and social media monitoring systems.
@@
Hinduism is the world’s oldest practising religion [18] with a population of more than 1.1 billion followers worldwide, mostly residing in India. Hinduism encompasses some of the oldest philosophical (Upanishads and Bhagavad Gita) [19, 20], epic and literary texts (Ramayana and Mahabharata) [21, 22]. Hinduism has contributed to modern science and technology with the invention of the binary and decimal number system [23] (also known as Hindu and Arabic numerals) [24] and systematic approach to scientific investigation using the philosophy of pramana (evidence) [25, 26]. Hinduism is based on non-violent principles that promote vegetarianism [27], with India having the largest vegetarian population in the world, with the majority of Hindus being partially vegetarian (specific days of vegetarian diet) [27]. Hinduism urges taking a violent stand in self-defence [28] and the fight for dharma (ethics) as promoted in the Bhagavad Gita [29]. The term Hinduphobia refers to hate speech and racial and cultural slurs about Hindu religious and cultural practices [13]. Although limited, a few studies have investigated the nature of Hinduphobia in television and popular culture, which became more prominent with the rise of the Internet and social media [14].
@@
Hinduism in India withstood its grounds with a history of violence and discrimination, given 800 years of foreign rule, with invasions leading to temple desecration [30]. Thousands of Hindu temples have been destroyed, desecrated, and converted into mosques and churches in history due to foreign rule with religious and political intervention [31]. This has been continuing with Hinduphobia in social media and political discourses, especially in countries where Hindus have been a minority, such as Pakistan and Bangladesh [32, 33]. In a majority-Hindu India, anti-Hindu political rhetoric led to the genocide and exodus of Hindus [34] from the Kashmir valley region in the 1980s [35]. Therefore, Hinduphobia is not a recent phenomenon; it has more than a thousand years of history, given the invasions and foreign rule of India [36]. Furthermore, Hinduphobic rhetoric has been one of the reasons for the forced conversion of Hindu women in Pakistan [37] and gradual changes in ownership and conversion of Hindu temples [38]. This has been an issue not just in India but in countries with a Hindu history, such as Indonesia [39].
@@
The Kumbh Mela, which is a major Hindu festival in India, was associated with a high number of COVID-19 cases during the Delta variant outbreak in April 2022 [40, 41, 42]. This association, driven by a political media narrative, led to negative portrayals of the Hindu community on social media platforms, including Twitter, Reddit, Instagram, and Facebook. These negative portrayals resulted in trends with hashtags such as #HinduVirus and #CoronaHindus, contributing to Hindubhobic statements and hate speech against the Hindu community. Previous studies have examined Hinduphobia in the context of historical events, political dynamics, academic biases, interreligious relations, diaspora experiences, and cultural practices [43]. Furthermore, there has been an upsurge in hate crimes against Hindus, as demonstrated by the cases of Khaniya Lal, Umesh Kohle, and Nupur Sharma [44]. These incidents highlight the escalating hostility towards Hindus during the pandemic and the need for religious tolerance in society. However, despite these concerning trends, no comprehensive computational framework has been developed to systematically monitor, detect, and analyze Hinduphobic content at scale during global crises. The COVID-19 pandemic has exacerbated communal tensions in India, which motivates our study to systematically study the rise of Hinduphobic sentiments connected to COVID-19 in India and the rest of the world.
@@
In recent years, LLMs have emerged as powerful tools for analysing and understanding public sentiment on social media. LLMs, such as GPT-3, PaLM, and GPT-4o [45, 46] can process vast amounts of textual data, revealing underlying patterns, sentiments, and discourse dynamics that shape public perception and societal narratives [47]. These models can be based on pre-training from a large data corpus and achieve strong performance accuracy in tasks such as decision-making, predictions, text generation, and text summarizations [48]. LLMs can detect nuanced language patterns and emerging trends, enabling researchers and policymakers to respond more effectively to rising social and political tensions [49], taking into account security and privacy [48]. Addressing the root causes of discrimination and fostering social cohesion is crucial; however, applying LLMs in this context has a lot of challenges. Data privacy, ethical considerations, and the potential for algorithmic bias must be carefully managed to ensure the deployment of LLMs for societal well-being [50]. LLMs must limit biases and maintain fairness, particularly when dealing with sensitive topics such as communal sentiments and hate speech [51].
@@
We present a computational framework for detecting and analyzing Hinduphobia during COVID-19. Our study introduces the ”Hinduphobic COVID-19 X (Twitter) Dataset” containing 8,000 annotated tweets collected from November 2019 to December 2022, employing rigorous human-in-the-loop annotation methodology [56]. The dataset categorizes tweets as Hinduphobic (negative) or positive/neutral, with all content related to Hinduism. Details are provided in (Section III-A1). The dataset is publicly available via Kaggle [^1] https://www.kaggle.com/datasets/ashutoshsingh22102/hinduphobic-covid-19-x-twitter-dataset-india.
We develop the Hinduphobic BERT (HP-BERT) model using multi-stage fine-tuning. First training on our ”Hinduphobic COVID-19 X (Twitter) Dataset” for hate detection, achieving 94.72% accuracy and outperforming five transformer models by 27.51% to 36.38% over baseline applications and 5.71% to 10.49% over fine-tuned versions. Subsequently, we fine-tune the model on the ”SenWave Dataset” [57] to add sentiment analysis capabilities [58, 59, 60], enabling HP-BERT to both detect Hinduphobic content and analyze sentiment while handling Hinglish content [61]. We apply HP-BERT to the ”Global COVID-19 Twitter dataset” [62], analyzing 27.4 million tweets from six countries: Australia, Brazil, India, Indonesia, Japan, and the United Kingdom. Statistical analysis reveals moderate correlations (r = 0.312-0.428) between COVID-19 case increases and Hinduphobic content volume. We analyze sentiment polarity scores to examine geographical and temporal variations in Hinduphobic discourse.
@@
The rest of the paper is organised as follows: Section II reviews the literature relevant to this study. Section III describes the datasets used, including the data collection and preprocessing processes, along with details of the training and testing of our BERT (HP-BERT) model. Section IV introduces the framework for our approach, and Section V presents the results. Finally, Section VI provides a comprehensive discussion, and Section VII concludes the paper with future research directions.
II Related Work
II-A Sentiment and Semantic Analysis
Sentiment and semantic analysis have become crucial components of NLP applications, particularly for analysing vast amounts of unstructured textual data from social media platforms [63, 64]. These methods involve understanding the emotional tone (sentiment) and the meaning (semantics) of textual content to extract insights regarding public opinion, trends, and biases [65]. LLMs have significantly improved the accuracy and scalability of these tasks by leveraging massive datasets and deep model architectures. LLMs can capture complex linguistic patterns and contextual nuances previously challenging to NLP. One of the earliest LLMs, the Bidirectional Encoder Representations from Transformers (BERT) model [66], has been widely adopted for sentiment analysis tasks [67]. Several customized BERT models have been developed, including news-recommendator BERT [68], fake news BERT [69], legal-BERT [70, 71], and medical-BERT [72] for respective domains. The multilingual-BERT (mBERT) [73] has been pre-trained on monolingual corpora in 104 languages, which has been designed to detect hate speech across multiple languages. DistilBERT [74] is a variant of BERT optimized for resource-limited applications. Moreover, advanced variants such as RoBERTa [75] have proven effective for nuanced hate speech detection across diverse settings. In the case of the detection of abuse and violence in NLP applications, BERT has been designed for various contexts and categories of hate speech. HateBERT [76] is trained on RAL-E, a large-scale dataset of Reddit comments from communities banned for being offensive, abusive, and hateful. These examples of customized BERT models and their variants demonstrate the adaptability and power of LLMs to address the complexities of hate speech detection in specific contexts, showcasing their impact in real-world NLP applications.
Sentiment and semantic analysis have been used to study domains using LLMs, including social media and text analysis. Wang and Chandra [17] used BERT-based sentiment analysis and reported that surges in COVID-19 cases correlated with spikes in Sinophobic sentiments. Similarly, Chandra et al. [77] used BERT-based sentiment analysis to study trends in The Guardian newspaper coverage of COVID-19, revealing a predominance of negative sentiments. In another study, Chandra and Kulkarni [78] utilised semantic and sentiment analysis to evaluate the Bhagavad Gita Sanskrit-English translations, comparing verse-by-verse consistency across selected translators. In the field of machine translation, Wang et al. [79] evaluated Google Translate for its accuracy in translating Mandarin Chinese, applying sentiment and semantic analysis to assess translation precision. Similarly, Shukla et al. [80] analysed Sanskrit-to-English translations of the Bhagavad Gita, uncovering inconsistencies in translating philosophical and metaphorical terms by Google Translate. Furthermore, Chandra and Saini [81] modelled the United States 2020 presidential elections using BERT-based sentiment analysis, providing insights into sentiment trends on social media before the elections. Chandra and Krishna [82] compared LSTM and BERT models to track public sentiment during the rise of COVID-19 cases in India. These studies collectively highlight the role of BERT-based models and LLMs in analysing sentiments across social and news media, illustrating their versatility and potential for understanding public opinion.
II-B Abuse and Misinformation
@@
The study of online activity drew significant attention from the research community due to the unique set of challenges faced during the COVID-19 pandemic, with concerns such as anti-vaccination and misinformation [83, 84, 85]. The substantially increased social media activity provided researchers with a wealth of data to analyze public sentiment and discourse. This surge in online interactions also exacerbated stress in society [86, 87] as individuals grappled with the rapid spread of misinformation, social isolation, and heightened fears about the pandemic. Chen et al. [88] curated a multilingual coronavirus Twitter dataset with over 123 million tweets for analysing the pandemic’s online discourse, reflecting major events and their impacts. Tahmasbi et al. [89] conducted one of the first studies on Sinophobic behaviour during the pandemic, analysing over 222 million tweets to reveal the rise in anti-Chinese sentiment.
Figure 1: Data collection and processing workflow from initial data extraction via Twitter API to manually labelling data points. This is followed by automatic labelling of the remaining tweets using GPT-3.5 Turbo using a human-in-the-loop strategy [56]. We removed incorrectly labeled tweets identified during verification to ensure final dataset quality.
@@
Cho et al. [90] focused on privacy concerns related to COVID-19 contact tracing apps such as TraceTogether, exploring cryptographic protocols and privacy-preserving techniques. The study highlighted the balance between privacy protection and public health benefits, emphasising effective anonymisation and additional privacy measures. Ferrara et al. [91] investigated the influence of X (Twitter) bots during the early COVID-19 outbreak, particularly in spreading political conspiracies and misinformation. Their curated dataset revealed how bots amplified divisive content during global emergencies. Kouzy et al. [92] examined COVID-19 misinformation on Twitter, identifying its rapid spread and sources. They found that misinformation often originated from informal groups rather than verified accounts, underscoring the need for robust monitoring and fact-checking mechanisms. Tahmasbi et al. [89] analysed Sinophobic content on social media platforms post-COVID-19, using word2vec models to trace the evolution of anti-Chinese sentiment. They highlighted the exacerbation of racial hostility during the pandemic across different online forums.
@@
Despite the extensive research on various aspects of the COVID-19 pandemic, there remains a gap in quantitative analysis regarding Hinduphobic sentiments during this period. However, no computational framework has been developed to systematically analyze Hinduphobia during global crises. This paper aims to fill this gap by examining the prevalence and impact of Hinduphobic discourse on social media platforms amidst the COVID-19 pandemic.
III Methods
III-A Datasets